Nitrogen and hydrogen gas react to form ammonia according to the reaction: N2(g)+3H2(g)→2NH3(g)


What is the limiting reactant?

Express your answer as a chemical formula.


Which reactant is in excess?

Express your answer as a chemical formula




1
Expert's answer
2021-05-31T02:25:36-0400

N2(g)+3H2(g)→2NH3(g)

so n(N2)=1mol

n (H2)=1mol

And we find amount compounds after the reaction:

n(N2)=1/coefficient=1/1=1mol

n(H2)=1/coefficient=1/3=0.333mol

So limit is H, so n (limit)=0.333mol

And excess is the N2 ,

so n(excess)=n(N2)-n(limit)=

=1-0.333=0.667mol